A population consistent of value 3,5,7 and 9





1)List all the possible of size n=3 when drawing with replacement and compute the mean of each sample





2) construct a sampling distribution of the mean





3) compute for the mean and variance of the of the sample means

1
Expert's answer
2022-03-03T17:58:36-0500

We have population values "3,5,7,9" population size "N=4" and sample size "n=3."

Thus, the number of possible samples which can be drawn with replacement is "4^3=64."


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Sample\ values & Sample\ mean(\bar{X}) \\ \hline
 3,3,3 & 3\\
 \hdashline
 3,3,5 & 11/3\\
 \hdashline
 3,3,7 & 13/3\\
 \hdashline
 3,3,9 & 5\\
 \hdashline
 3,5,3 & 11/3\\
 \hdashline
 3,5,5 & 13/3\\
 \hdashline
 3,5,7 & 5\\
 \hdashline
 3,5,9 & 17/3\\
 \hdashline
 3,7,3 & 13/3\\
 \hdashline
 3,7,5 & 5\\
 \hdashline
 3,7,7 & 17/3\\
 \hdashline
 3,7,9 & 19/3\\
 \hdashline
 3,9,3 & 5\\
 \hdashline
 3,9,5 & 17/3\\
 \hdashline
 3,9,7 & 19/3\\
 \hdashline
 3,9,9 & 7\\
 \hdashline
 5,3,3 & 11/3\\
 \hdashline
 5,3,5 & 13/3\\
 \hdashline
 5,3,7 & 5\\
 \hdashline
 5,3,9 & 17/3\\
 \hdashline
 5,5,3 & 13/3\\
 \hdashline
 5,5,5 & 5\\
 \hdashline
 5,5,7 & 17/3\\
 \hdashline
 5,5,9 & 19/3\\
 \hdashline
 5,7,3 & 5\\
 \hdashline
 5,7,5 & 17/3\\
 \hdashline
 5,7,7 & 19/3\\
 \hdashline
 5,7,9 & 7\\
 \hdashline
 5,9,3 & 17/3\\
 \hdashline
 5,9,5 & 19/3\\
 \hdashline
 5,9,7 & 7\\
 \hdashline
 5,9,9 & 23/3\\
 \hdashline
 7,3,3 & 13/3\\
 \hdashline
 7,3,5 & 5\\
 \hdashline
 7,3,7 & 17/3\\
 \hdashline
 7,3,9 & 19/3\\
 \hdashline
 7,5,3 & 5\\
 \hdashline
 7,5,5 & 17/3\\
 \hdashline
 7,5,7 & 19/3\\
 \hdashline
 7,5,9 & 7\\
 \hdashline
 7,7,3 & 17/3\\
 \hdashline
 7,7,5 & 19/3\\
 \hdashline
 7,7,7 & 7\\
 \hdashline
 7,7,9 & 23/3\\
 \hdashline
 7,9,3 & 19/3\\
 \hdashline
 7,9,5 & 7\\
 \hdashline
 7,9,7 & 23/3\\
 \hdashline
 7,9,9 & 25/3\\
 \hdashline
 9,3,3 & 5\\
 \hdashline
 9,3,5 & 17/3\\
 \hdashline
 9,3,7 & 19/3\\
 \hdashline
 9,3,9 & 7\\
 \hdashline
 9,5,3 & 17/3\\
 \hdashline
 9,5,5 & 19/3\\
 \hdashline
 9,5,7 & 7\\
 \hdashline
 9,5,9 & 23/3\\
 \hdashline
 9,7,3 & 19/3\\
 \hdashline
 9,7,5 & 7\\
 \hdashline
 9,7,7 & 23/3\\
 \hdashline
 9,7,9 & 25/3\\
 \hdashline
 9,9,3 & 7\\
 \hdashline
 9,9,5 & 23/3\\
 \hdashline
 9,9,7 & 25/3\\
 \hdashline
 9,9,9 & 9\\
 \hdashline
\end{array}"



2) The sampling distribution of the sample mean "\bar{X}" is


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 & \bar{X} & f & f(\bar{X}) & Xf(\bar{X})& X^2f(\bar{X}) \\ \hline
 & 3 & 1 & 1/64 & 3/64 & 27/192\\
 \hdashline
 & 11/3 & 3 & 3/64 & 11/64 & 121/192 \\
 \hdashline
 & 13/3 & 6 & 6/64 & 26/64 & 338/192\\
 \hdashline
 & 5 & 10 & 10/64 & 50/64 & 750/192 \\
 \hdashline
 & 17/3 & 12 & 12/64 & 68/64 & 1156/192 \\
 \hdashline
 & 19/3 & 12 & 12/64 & 76/64 & 1444/192 \\
 \hdashline
 & 7 & 10 & 10/64 & 70/64 & 1470/192 \\
 \hdashline
 & 23/3 & 6 & 6/64 & 46/64 & 1058/192 \\
 \hdashline
 & 25/3 & 3 & 3/64 & 25/64 & 625/192 \\
 \hdashline
 & 9 & 1 & 1/16 & 9/64 & 243/192 \\
 \hdashline
 Sum= & & 64 & 1 & 6 & 113/3\\
 \hdashline
\end{array}"

"\mu=\dfrac{3+5+7+9}{4}=6"


"\sigma^2=\dfrac{1}{4}((3-6)^2+(5-6)^2+(7-6)^2+(9-6)^2)"

"=5"


The mean of the sample means is


"\mu_{\bar{X}}=E(\bar{X})=6"




"Var(\bar{X})=\sigma^2_{\bar{X}}=E(\bar{X}^2)-(E(\bar{X}))^2"


"=\dfrac{113}{3}-(6)^2=\dfrac{5}{3}"

"\mu_{\bar{X}}=\mu"

"\sigma_{\bar{X}}^2=\dfrac{\sigma^2}{n}"